OpenStudy (anonymous):

Find the distance between the two points (5,-3) and (-1,-6) and the midpoint of the line segment that joins the two points.

OpenStudy (anonymous):

\[y _{2}-y_{1}/x _{2}-x_{1}\]

OpenStudy (anonymous):

plug your points into that question

OpenStudy (anonymous):

that is your slope matt.... distance would be \[\sqrt{(x _2-x _1)^2+(y _2-y _1)^2}\]

OpenStudy (anonymous):

mid point will be \[(\frac{x _1+x _2}{2},\frac{y _1+y _2}{2})\]

OpenStudy (anonymous):

thank you very much.

OpenStudy (anonymous):

so your midpont would be \[\frac{5+-1}{2},\frac{-3+-6}{2}= (2,\frac{-9}{2})\]

OpenStudy (anonymous):

distance will be \[\sqrt{(-1-5)^2+(-6-(-3))^2}=\sqrt{(-6)^2+(-3)^2}=\sqrt{36+9}=\sqrt{45}=\sqrt{9*5}=3\sqrt{5}\]

OpenStudy (anonymous):

\[\sqrt{45}=\sqrt{9*5}=3\sqrt{5}\]